1. Thirty percent of all resumes received by a corporation for a management position are from females. Ten resumes will be received tomorrow.

a. What is the probability that exactly 4 of the resumes will be from females?

b. What is the expected number of resumes from women?

c. What is the variance of the number of resumes from women?
Answer #1

Solution

P(Resume received from Female) = 0.3

No. Of sample = 10

Solution(a)

P(X=4) = 10C4*(0.3)^4*(0.7)^6 = 0.2001

Solution(b)

Expected number of resumes from women = No.of female * probability of resume received from female = 0.3*10 = 3

Solution(c)

Variance of the number of resumes from women = N*P*Q = 10*0.3*0.7 = 2.1